OpenStudy (anonymous):

Simplify the expression. 3 + 2 • (8 – 3) A. 25 B. 37 C. 16 D. 13 A?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

D @PizzaLover123

OpenStudy (anonymous):

How?:P I got A

OpenStudy (anonymous):

But it is D

OpenStudy (anonymous):

5 • 5 = 25

OpenStudy (anonymous):

(8-3)=

OpenStudy (anonymous):

(5)

OpenStudy (anonymous):

5*2=10

OpenStudy (anonymous):

3 + 2 • (5) 3 + 2 = 5 • (5) 5 • 5 = 25

OpenStudy (anonymous):

you must always start with the brackets

OpenStudy (anonymous):

3 + 2 • (8 – 3) (8-3)=5 5*2=10 10+3=13

OpenStudy (anonymous):

@PizzaLover123

OpenStudy (anonymous):

so it will be D=13

OpenStudy (anonymous):

ok @PizzaLover123 Do you get it

OpenStudy (anonymous):

????????????????????

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

can you Medal me or Fan me @PizzaLover123